Find the volume of the metal used in a hollow cylinder open at both ends whose external diameter is 28 cm, thickness 3.5 cm and height 10 cm.
Answer by lwsshak3(11628)

Find the volume of the metal used in a hollow cylinder open at both ends whose external diameter is 28 cm, thickness 3.5 cm and height 10 cm.
***
internal diameter=28-3.5=24.5 cm
volume= internal area of end*height
area of end=πd^2/4, d=diameter
Volume=π*d^2/4)*10
=π(24.5)^2/4*10=1500.625π
=4714.35 cm^3
..
external diameter=28 cm
volume= external area of end*height
area of end=πd^2/4, d=diameter
Volume=π*(d^2/4)*10
=π(28)^2/4*10=1960π
=6167.52 cm^3
..
6167.52-4714.35=1453.17 cm^3
volume of the metal used in a hollow cylinder=1453.17 cm^3